OpenStudy (anonymous):

What is the common difference between the elements of the arithmetic sequence below? –18, –22.5, –27, –31.5, –36

OpenStudy (anonymous):

it would be -4.5. you can get that substracting each element from the next one. -22.5-(-18)=-22.5+18=-4.5

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

can you help me with another one? you don't mind